Circular No. 6295

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CASSIOPEIAE 1995 R. Gonzalez-Riestra, IUE/Vilspa; S. N. Shore, Indiana University at

South Bend; S. Starrfield, Arizona State University; and J. Krautter, Landessternwarte Heidelberg, report: “Nova Cas 1995 has been monitored with IUE since about one month after discovery. Integrated fluxes in the 120– 200-nm range are (in units of 10^{−11} erg cm^{−2} s^{−1}): 1995 Sept. 22.1 UT, 2.90; Dec. 22.8, 1.30; 23.8, 1.80; 1996 Jan. 6.8, 9.61; 17.3, 15.2. The initial spectrum showed a sloping continuum increasing toward longer wavelengths characteristic of the fireball phase. The spectrum obtained during the optical flare in Dec. 1995 (IAUC 6275, 6283) showed a dramatic increase in the iron peak absorption shortward of 170 nm, and windows in the continuum developed as the flux increased. The reddening, as estimated from the strength of the 220-nm absorption feature, is E(B − V) = 0.6. This value is consistent with an evolution at constant bolometric luminosity. Highresolution (R = 10 000) spectra obtained on Jan. 6 show a weak P-Cyg emission feature at Mg II 280 nm with broad absorption extending to −500 km/s. The line profile is not consistent with the presence of a wind and rather resembles the profiles seen in the optically-thick stages of other novae. The mean radial velocity for the interstellar component is −15 km/s. The N III] 175-nm and C III] 191-nm lines appeared in the Jan. 6 spectrum, characteristic of the prenebular phase. No other unambiguous emission has been detected in the short-wavelength spectra. The absorption spectrum is still very strong.”

RX J1239.5+2108 = S 10932 COMAE BERENICES G. A. Richter and P. Kroll, Sonneberg and Tautenburg Observatories;

and J. Greiner, Max Planck Institute for Extraterrestrial Physics, report: “The recent reports via the VSNET exploder of an outburst of S 10932 Com, a cataclysmic variable at α = 12^{h}39^{m}32^{s}.1, δ = +21^{o}08′06′′ (equinox 2000.0; Richter and Greiner 1995, in Cataclysmic Variables, ed. A. Bianchini et al., ASSL 205, p. 177) motivated us to improve the eclipse lightcurve (Wenzel et al. 1995, IBVS 4182). Using CCD photometry with the Sonneberg 0.60-m Cassegrain telescope during 1994 May 11–1996 Jan. 4 and archival photographic eclipse observations since 1973, we derive for the minimum the ephemeris HJD = 2449486.48166 + 0.0870386727E. The many observed eclipses are of varying depth (1.5–4.4 mag in the R band). The maximum depth occurred on 1996 Jan. 4–5, i.e., a few days after the recent outburst. This outburst is the third after those in 1959 and 1988, implying a possible cycle length of 8 years.”

1996 January 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden